Your aunt made you a beneficiary of her insurance policy. It is worth $200,000 and it will earn 5% interest over your expected lifetime of 45 years. How much income could you get if you take the income beginning immediately ( in other words the insurance is converted to an annuity due)?
Answer #1

PV of Annuity Due = Annuity + Annuity * [{1 - (1 + r)-(n-1)} / r]

$200,000 = C + C*[{1 - (1 + 0.05)-(45-1)} / 0.05]

$200,000 = C + C*[0.8831 / 0.05]

$200,000 = C + C*[17.6628]

$200,000 = C*[18.6628]

C = $200,000 / 18.6628 = $10,716.52